# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ Abstract
Nonreciprocal interactions that violate Newton’s law ’actio=reactio’ are ubiquitous in nature and are currently intensively investigated in active matter, chemical reaction networks, population dynamics, and many other fields. An outstanding challenge is the thermodynamically consistent formulation of the underlying stochastic dynamics that obeys local detailed balance and allows for a rigorous analysis of the stochastic thermodynamics of non-reciprocally interacting particles. Here, we present such a framework for a broad class of active systems and derive by systematic coarse-graining exact expressions for the macroscopic entropy production. Four independent contributions to the thermodynamic dissipation can be identified, among which the energy flux sustaining vorticity currents manifests the presence of non-reciprocal interactions. Then, Onsager’s non-reciprocal relations, the fluctuation-response relation, the fluctuation relation and the thermodynamic uncertainty relations for non-reciprocal systems are derived. Finally, we demonstrate that our general framework is applicable to a plethora of active matter systems and chemical reaction networks and opens new paths to understand the stochastic thermodynamics of non-reciprocally interacting many-body systems.
@@ 1 Introduction
Figure 1: Scheme of various levels of coarse-graining considered in this paper, microscopic, mesoscopic, macroscopic-fluctuating, macroscopic-deterministic and hydrodynamic. (a) Microscopic particles are confined to the lattice site (denoted by the black ##\##). Four different types of particles are illustrated: triangle (blue), square (pink), pentagon (green) and, circle (orange). (b/c) The mesoscopic/macroscopic description is obtained for the fluctuating particle number/density. (d) The straight interface between the density fields represents the suppression of the fluctuation and convergence to the mean-field dynamical description. (e) Cyan and Yellow denote the relevant hydrodynamic order parameter that does not necessarily preserve the microscopic properties.
@@
The discovery of fluctuation relations shed new light on the classical field of thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and established the refined framework of stochastic thermodynamics (ST) [20, 21]. Stochastic thermodynamics accomplishes to define the first and second laws of thermodynamics for a single stochastic transition  [^1] We stick to the convention of using a stochastic transition instead of a stochastic trajectory usually used. This is because a stochastic trajectory is visualized as a collection of sequential stochastic transitions. The Markovian property of the transitions ensures this equivalence, provided that the initial probability distributions are the same. We clarify this notion to avoid confusion with the trajectory-based approach used to compute the information-theoretic definition of entropy production. The information-theoretic approach is not thermodynamically consistent if each microscopic transition is not properly resolved or the transition rates do not satisfy the local detailed balance condition. This is achieved by defining the energy, work, heat and, entropy production for a stochastic transition, such that the thermodynamic quantities have probability distributions. Fluctuations vanish in the thermodynamic limit, hence the resulting stochastic thermodynamic quantities are peaked to the deterministic thermodynamic description. Recovering the traditional thermodynamic description in the thermodynamic limit. Stochastic thermodynamics applies to small systems with inherent stochastic fluctuations in the presence of an externally imposed driving of control parameters. The validity of this framework relies on a key constraint of time-scale separation which results in Markovian property for the transitions: states of the stochastic process comprise the slow degrees of freedom and the collective fast degrees of freedom form the thermal environment (bath). The Local Detailed Balance (LDB) condition establishes the equivalence between the forward and backward transition rates and the thermodynamic transition cost supported by the environment [23, 24] and ensures thermodynamic consistency for all transitions.
@@
Newton’s third law of motion states that for every action force, there is equal and opposite reaction force. Multiple real-world counter-examples to Newton’s third law of motion exist. The systems violating Newton’s third law of motion are called non-reciprocal [25]. In particular, a multitude of biological systems are frontier to study and understand the implications of the violation of Newton’s third law of motion. For instance, toy models for ecological systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], stochastic systems [37, 38, 39], neural networks [40, 41, 42, 43, 44, 45, 46], a mixture of non-reciprocal particles [47], birds with vision cones [48], solids with odd elasticity [49], particles breaking action-reaction symmetry [50, 51], chiral active particles [52, 53], non-reciprocal flocking models [54, 55], non-reciprocal frustration [56], systems described by non-reciprocal Cahn-Hillard equations [57, 58, 59, 60, 61, 62, 63]. Recent developments in understanding the microscopic dynamics of non-reciprocal systems have revealed more sophisticated behaviors and phases that are otherwise absent in reciprocal systems. For example, travelling waves, oscillations, spiral waves, and effervescence [54, 55, 57, 58, 59, 60, 61, 62, 63]. The non-reciprocal phases have been understood by delineating the underlying topological properties [54].
@@
Active matter and non-reciprocal systems are conventionally analyzed with the hydrodynamic description [64, 65], which successfully recovers the phases observed for the microscopic dynamics [55]. Deriving hydrodynamic equations of motion requires identifying the relevant degrees of freedom termed order parameter with the help of symmetry arguments: a top-down approach [64, 65]. Subsequently, one can incorporate the order parameter fluctuations by using more sophisticated exact coarse-graining techniques [66, 67]. This procedure is usually effective and reliable in understanding the system dynamics, phases, and phase transitions.
@@
However, for microscopic systems, the hydrodynamic description fails quantitatively and qualitatively in the small particle number regime. For example, the onset of the phase transition predicted by the hydrodynamic description is different from that predicted by the microscopic description [68]. Moreover, it fails to predict the order of the phase transition which is usually addressed by introducing ad-hoc noise corrections [68]. The hydrodynamic description also fails to adequately describe and understand the thermodynamic implications at the macro/meso-scale because of coarse-graining. The effective hydrodynamic description does not resolve the microscopic dissipation at the hydrodynamic scale. In particular, the microscopic states with different thermodynamic properties are coarse-grained to a single-order parameter. In addition, the mapping between the microscopic and macroscopic/mesoscopic control parameters is less transparent. We aim to construct the macrostate dynamics using a bottom-up approach. We want to implement the coarse-graining procedure only up to an order in which the microscopic dissipation is both qualitatively and quantitatively preserved, and therefore avoid a subsequent coarse-graining to obtain an effective description in terms of relevant order parameters.
@@
The coarse-graining scheme is shown in fig. 1. Microscopic particles with different dynamic and thermodynamic microscopic properties exist. Lumping together microstates with identical thermodynamic and physical properties forms a mesostate: the particle number. The macroscopic limit corresponds to taking the limit of a large number of particles per site. It defines the scaling from the intensive density order parameter to the extensive particle number. The densities are fluctuating stochastic fields, denoted by the rough edges. The thermodynamic limit suppresses the macroscopic fluctuations that lead to the reaction-diffusion dynamics for deterministic density fields. In comparison, the top-down construction of an effective hydrodynamic description fails to track the microscopic thermodynamic dissipation qualitatively and quantitatively.
@@
Despite much progress in understanding the dynamics of active matter on different scales, its implications for the thermodynamics of non-reciprocal systems remain elusive. For reciprocal active matter systems the stochastic thermodynamics has been formulated for active Brownian particles [69, 70, 71] and for flocking models [72] and fluctuation relations have been derived [73, 74]. To quantify thermodynamic irreversibility, an information-theoretic measure has been utilized, as for instance for the Active Ising model [75] and the non-reciprocal Cahn-Hillard equation [76, 77, 78], which has been proven to be a good statistical measure to compute lower bounds for the thermodynamic entropy production [79, 80, 81, 82, 83, 84, 85]. However, the information-theoretic definition of irreversibility does not necessarily coincide with the thermodynamic entropy production [79, 80, 81, 82, 83, 84, 85]. The information theoretic irreversibility measure can be defined without the state transitions satisfying local detailed balance (LDB), thus lacking thermodynamic consistency [75, 76, 77].
@@
We therefore combine in this paper the three main motifs mentioned so far: Stochastic Thermodynamics and Coarse-graining for the Nonreciprocal Interacting Particles. We aim to formulate a microscopic Markov jump process description for the interacting non-reciprocal particles, which enables us to define their dynamics and thermodynamics in a thermodynamically consistent way. Subsequently, we aim to implement a coarse-grained mesoscopic/macroscopic description that preserves the microscopic thermodynamic dissipation. We ensure this by properly identifying the LDB condition on the mesoscale/macroscale. We obtain the thermodynamically consistent dynamics for the time evolution of the mesostate/macrostate. Importantly, this ensures the thermodynamics for the non-reciprocal systems is defined at the mesoscale/macroscale. This allows us to establish the connection between the microscopic and macroscopic/mesoscopic control parameters, which allows a qualitative and quantitative equivalence between the dynamics and thermodynamic physical quantities. Importantly, this extends the applicability of Stochastic Thermodynamics to coarse-grained many-body descriptions.
@@
We define the class of thermodynamically consistent lattice gas models for non-reciprocally interacting microscopic particles coupled to an external driving reservoir [1, 13, 20]. LDB for the microscopic system ensures the thermodynamic consistency of these models. This allows us to define the microscopic EPR for the externally driven non-reciprocal particles using the Markov Jump Process (MJP) and its dynamics governed by the Master equation [1, 13, 20]. Further, we define the coarse-grained state description (mesoscopic/macroscopic) by lumping microstates with identical dynamical and thermodynamical properties [86]. The mesoscopic/macroscopic scale is defined for the particle number/density as a coarse-grained state. The coarse-grained description is chosen flexibly depending on the system-specific requirement. We employ the Doi-Peliti exact coarse-graining method that preserves microscopic discreetness at the coarse-grained description [86].
@@
Next, we formulate the transitions between the coarse-grained state using the LDB in a thermodynamically consistent way. This enables us to quantify and decompose the coarse-grained state dissipation using ST for any non-reciprocal system. In addition, the dynamic and thermodynamic equivalence of the microscopic and coarse-grained systems is ensured. The fluctuating dynamics of the coarse-grained states generalize the Macroscopic Fluctuating Theory (MFT) for the non-reciprocal systems [87]. We obtain the non-quadratic dissipation function that plays an important role in the exact quantification of the EPR (particularly for the far-from-equilibrium systems [88]. We decompose the EPR into its linearly independent contributions, namely orthogonal decomposition. It consists of four contributions, relaxation of the reciprocal interaction energy functional, non-reciprocal interaction, coupling to the external reservoir, and external driving work.
@@
Finally, we demonstrate the non-reciprocal and non-equilibrium analogues of the reciprocal systems. For instance, the irreversible thermodynamics relations: the Onsager’s non-reciprocal relations, fluctuation-response relations, higher-order non-linear responses and, the stochastic thermodynamic relations: fluctuation relations, thermodynamic kinetic uncertainty relation. In the deterministic macroscopic limit, we recover the irreversible thermodynamics, the Arrhenius’s transition state theory (LDB) and reaction-diffusion systems (dynamics) for the interacting non-reciprocal density fields [89, 64, 65]. We elaborate on the application of our framework using a few prototypical models. Our framework extends the applicability of stochastic thermodynamics for non-reciprocal and externally driven systems across different observation scales.
@@ 2.1 Microscopic Boltzmann weight
We consider a multi-particle system that comprises several particle types and lives in discretized space, a regular lattice with lattice constant $l$, continuous time, and periodic boundary conditions. Physically, the lattice constant corresponds to using the microscopic diffusive length scale of the particles as a space measurement unit. The continuum limit $l\to 0$ leads to oversimplified coarse-grained description that does not account for the inherent diffusive length-scale of the system. Therefore, we consider $l=1$ in this paper unless we specify the continuum space limit $l\to 0$. The number of particles of type $i$ at the lattice site index ##\## is denoted by $N_{i}^{\#}$. The configuration of the lattice is denoted by $\{N\}$. The dimension of the $\{N\}$ is decided by the lattice size and the number of particle types. The volume of the d-dimensional lattice space is denoted by $\mathcal{V}$. The faster degrees of freedom forms the environment which supports the thermodynamic cost for all microstate changes of the system which consist of the slower degrees of freedom.
The microscopic interaction energy experienced by type $i$ particle due to the type $j$ particle is denoted by $v_{ij}$. $v_{ij}<0$ implies type $i$ particle are attracted towards type $j$ particles, $v_{ij}>0$ holds true for the repulsive interaction. Formulated analogously, $v_{ij}$ is the thermodynamic cost supported by the environment to place a type $i$ particle in the presence of a type $j$ particle. The violation of the actio=reactio symmetry implies $v_{ij}\neq v_{ji}$ microscopically. $v_{ij}=v_{ji}$ implies the particles are reciprocal and respect the actio=reactio symmetry. To differentiate and quantify the actio=reactio symmetry preserving and breaking terms, we define the reciprocal $v_{ij}^{r}=(v_{ij}+v_{ji})/2$ and non-reciprocal $v_{ij}^{nr}=(v_{ij}-v_{ji})/2$ interaction energy between the type $i$ and type $j$ particles. $v_{ij}^{r}$ and $v_{ij}^{nr}$ correspond to the actio=reactio symmetry preserving and breaking components respectively. Importantly, this fixes ‘the orthogonal reference gauge’ such that the symmetry $v_{ij}^{r}=v_{ji}^{r}$ and anti-symmetry $v_{ij}^{nr}=-v_{ji}^{nr}$ is satisfied by construction.
The reciprocal and non-reciprocal components of the microscopic Boltzmann weight $\epsilon_{i}^{\#}$ of the type $i$ particle at ##\## are defined as $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ respectively  [^2] We drop the lattice index ##\## for $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ for the brevity. However, it is assumed that these objects will be evaluated with the required lattice ##\## and type $i$ indexes., such that $\epsilon_{i}^{\#}=\epsilon_{i}^{r}+f_{i}^{nr}$,

### v3

@@ Abstract
Nonreciprocal interactions that violate Newton’s law ‘actio=reactio’ are ubiquitous in nature and are currently intensively investigated in active matter, chemical reaction networks, population dynamics, and many other fields. An outstanding challenge is the thermodynamically consistent formulation of the underlying stochastic dynamics that obeys local detailed balance and allows for a rigorous analysis of the stochastic thermodynamics of non-reciprocally interacting particles. Here, we present such a framework for a broad class of active systems and derive by systematic coarse-graining exact expressions for the macroscopic entropy production. Four independent contributions to the thermodynamic dissipation can be identified, among which the energy flux sustaining vorticity currents manifests the presence of non-reciprocal interactions. Then, Onsager’s non-reciprocal relations, the fluctuation-response relation, the fluctuation relation and the thermodynamic uncertainty relations for non-reciprocal systems are derived. Finally, we demonstrate that our general framework is applicable to a plethora of active matter systems and chemical reaction networks and opens new paths to understand the stochastic thermodynamics of non-reciprocally interacting many-body systems.
@@ 1 Introduction
Figure 1: Scheme of various levels of coarse-graining considered in this paper, microscopic, mesoscopic, macroscopic-fluctuating, macroscopic-deterministic and hydrodynamic. (a) Microscopic particles are confined to the lattice site (denoted by the black #\#). Four different types of particles are illustrated: triangle (blue), square (pink), pentagon (green) and, circle (orange). (b/c) The mesoscopic/macroscopic description is obtained for the fluctuating particle number/density. (d) The straight interface between the density fields represents the suppression of the fluctuation and convergence to the mean-field dynamical description. (e) Cyan and Yellow denote the relevant hydrodynamic order parameter that does not necessarily preserve the microscopic properties.
@@
The discovery of fluctuation relations has shed new light on the classical field of thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and established the refined framework of stochastic thermodynamics (ST) [20, 21]. Stochastic thermodynamics provides a framework to define the first and second laws of thermodynamics for a single stochastic transition  [^1] We adhere to the convention of using a stochastic transition instead of a stochastic trajectory, since a stochastic trajectory can be visualized as a collection of sequential stochastic transitions. The Markovian property of the transitions ensures this equivalence, provided that the initial probability distributions (boundary terms) are the same. We clarify this notion to avoid confusion with the trajectory-based approach used to compute the information-theoretic definition of entropy production. The information-theoretic approach is not necessarily thermodynamically consistent if each microscopic transition is not properly resolved or if the transition rates do not satisfy the local detailed balance condition.. This is achieved by defining the energy, work, heat, and entropy production for a stochastic transition, so that these thermodynamic quantities have probability distributions. Fluctuations vanish in the thermodynamic limit, where the resulting stochastic thermodynamic quantities converge to the deterministic thermodynamic description, thereby recovering classical thermodynamics. Stochastic thermodynamics applies to small systems with inherent stochastic fluctuations in the presence of externally imposed driving of control parameters. The validity of this framework relies on a key constraint of time-scale separation, which results in the Markovian property for the transitions: the states of the stochastic process comprise the slow degrees of freedom, while the collective fast degrees of freedom form the thermal environment (bath). The Local Detailed Balance (LDB) condition establishes the equivalence between the forward and backward transition rates for the system and the thermodynamic transition cost supported by the environment [23, 24], ensuring thermodynamic consistency for all transitions.
@@
Newton’s third law of motion states that for every action force, there is an equal and opposite reaction force. However, multiple real-world counter-examples to Newton’s third law exist. Systems that violate Newton’s third law are called non-reciprocal [25]. In particular, a multitude of biological systems provides frontiers for studying and understanding the implications of such violations. For instance, toy models for ecological systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], stochastic systems [37, 38, 39], neural networks [40, 41, 42, 43, 44, 45, 46], mixtures of non-reciprocal particles [47], birds with vision cones [48], solids with odd elasticity [49], particles breaking action-reaction symmetry [50, 51], chiral active particles [52, 53], non-reciprocal flocking models [54, 55], non-reciprocal frustration [56], and systems described by non-reciprocal Cahn-Hilliard equations [57, 58, 59, 60, 61, 62, 63]. Recent developments in understanding the microscopic dynamics of non-reciprocal systems have revealed more sophisticated behaviors and phases absent in reciprocal systems, such as traveling waves, oscillations, spiral waves, and effervescence [54, 55, 57, 58, 59, 60, 61, 62, 63]. The non-reciprocal phases have been further elucidated by delineating the underlying topological properties [54].
@@
Active matter and non-reciprocal systems are conventionally analyzed using hydrodynamic descriptions [64, 65], which successfully recover the phases observed in the microscopic dynamics [55]. Deriving hydrodynamic equations of motion requires identifying the relevant degrees of freedom, termed order parameters, with the help of symmetry arguments: a top-down approach [64, 65]. Subsequently, fluctuations of the order parameters can be incorporated using more sophisticated exact coarse-graining techniques [66, 67]. This procedure is usually effective and reliable for understanding system dynamics, phases, and phase transitions.
@@
However, for microscopic systems, the hydrodynamic description fails both quantitatively and qualitatively in the small particle number regime. For example, the onset of the phase transition predicted by the hydrodynamic description differs from that predicted by the microscopic description [68]. Moreover, it does not correctly predict the order of the phase transition, which is usually addressed by introducing ad-hoc noise corrections [68]. The hydrodynamic description also fails to adequately describe and interpret the thermodynamic implications at the macro/meso-scale because of coarse-graining. The effective hydrodynamic description does not resolve microscopic dissipation at the hydrodynamic scale. In particular, microscopic states with different thermodynamic properties are coarse-grained into a single order parameter. In addition, the mapping between microscopic and macroscopic/mesoscopic control parameters is less transparent. We aim to construct the macrostate dynamics using a bottom-up approach, implementing the coarse-graining procedure only to an order at which the microscopic dissipation is both qualitatively and quantitatively preserved, thereby avoiding a subsequent coarse-graining to obtain an effective description in terms of relevant order parameters.
@@
The coarse-graining scheme is illustrated in fig. 1. Microscopic particles with different dynamic and thermodynamic properties exist. Lumping together microstates with identical thermodynamic and physical properties forms a mesostate: the particle number. The macroscopic limit corresponds to taking the limit of a large number of particles per site. It defines the scaling from the intensive density order parameter to the extensive particle number. The densities are fluctuating stochastic fields, indicated by the rough edges. The thermodynamic limit suppresses macroscopic fluctuations, leading to reaction-diffusion dynamics for deterministic density fields. In comparison, the top-down construction of an effective hydrodynamic description fails to track microscopic thermodynamic dissipation both qualitatively and quantitatively.
@@
Despite significant progress in understanding the dynamics of active matter on different scales, their implications for the thermodynamics of non-reciprocal systems remain elusive. For reciprocal active matter systems, stochastic thermodynamics has been formulated for active Brownian particles [69, 70, 71] and for flocking models [72], and fluctuation relations have been derived [73, 74]. To quantify thermodynamic irreversibility, information-theoretic measures have been utilized, for instance, for the Active Ising model [75] and the non-reciprocal Cahn-Hilliard equation [76, 77, 78, 79], which have been proven to be good statistical measures for computing lower bounds on thermodynamic entropy production [80, 81, 82, 83, 84, 85, 86]. However, the information-theoretic definition of irreversibility does not necessarily coincide with thermodynamic entropy production [80, 81, 82, 83, 84, 85, 86]. The information-theoretic measure can be defined without the state transitions satisfying local detailed balance (LDB), thus lacking thermodynamic consistency [75, 76, 77, 78].
@@
In this paper, we combine the three main motifs mentioned so far: Stochastic Thermodynamics, Coarse-graining, and Non-reciprocal Interacting Particles. We aim to formulate a microscopic Markov jump process description for interacting non-reciprocal particles, which enables a thermodynamically consistent definition of their dynamics and thermodynamics. Subsequently, we aim to obtain a coarse-grained mesoscopic/macroscopic description that preserves microscopic thermodynamic dissipation. We ensure this by properly identifying the LDB condition at the mesoscale/macroscale. This allows us to obtain thermodynamically consistent dynamics for the time evolution of the mesostate/macrostate, ensuring that thermodynamics for non-reciprocal systems is defined at these scales. This procedure establishes the connection between microscopic and macroscopic/mesoscopic control parameters, enabling qualitative and quantitative equivalence between dynamics and thermodynamic quantities. Importantly, this extends the applicability of stochastic thermodynamics to coarse-grained many-body descriptions.
Throughout this work, the definition of a ‘microscopic’ non-reciprocally interacting particle refers to the ‘minimal’ experimentally relevant observation scale, where an ‘effective’ non-reciprocity is observed. This non-reciprocity may arise from asymmetric interactions with the surrounding environment, complex internal dynamics, or the cognitive capabilities of living systems, corresponding to active matter models, chemical reaction networks, and population dynamics models, respectively. For example, the non-reciprocal attraction–repulsion dynamics of dogs and sheep. Although dogs and sheep are composed of microscopic atoms and molecules that may satisfy Newton’s third law at the smallest scales, a dog and a sheep are treated as ‘effective’ particles exhibiting non-reciprocal interactions due to their complex internal cognitive capabilities. We treat such an ‘effective non-reciprocal particle’ as the fundamental microscopic object of our study. The notion of a ‘microscopic particle’ refers to such an independent degree of freedom. Questions such as ‘How does such non-reciprocity originate, and what is the energetic cost to generate it?’ are not the aim of this work. Rather, we focus on the open question: given the abundant existence of ‘effective’ non-reciprocal particles, is it possible to formulate a thermodynamically consistent microscopic theoretical description of their dynamics? We focus on a broad class of lattice models with non-reciprocal interactions, for which we introduce a thermodynamically consistent dynamics. Then it is a well posed question to derive the stochastic thermodynamics for it, but we intended this as a purely theoretical framework without any concrete experimental realization in mind.
@@
We define a class of thermodynamically consistent lattice gas models for non-reciprocally interacting microscopic particles coupled to an external driving reservoir [1, 14, 20]. We achieve this by formulating the LDB condition for the dynamics of microscopic non-reciprocal particles. The LDB for the microscopic system ensures thermodynamic consistency of these models. This allows us to define the microscopic entropy production rate (EPR) for externally driven non-reciprocal particles using a Markov Jump Process (MJP) with dynamics governed by the Master equation [1, 14, 20]. Further, we define a coarse-grained description (mesoscopic/macroscopic) by lumping microstates with identical dynamical and thermodynamic properties [87]. The mesoscopic/macroscopic scale is defined for particle number/density as a coarse-grained state. The coarse-grained description is chosen flexibly depending on system-specific requirements. We employ the Doi-Peliti exact coarse-graining method, which preserves microscopic discreteness in the coarse-grained description [87].
@@
Next, we formulate transitions between coarse-grained states using the LDB in a thermodynamically consistent way. This enables us to quantify and decompose coarse-grained state dissipation using stochastic thermodynamics for any non-reciprocal system. In addition, the dynamic and thermodynamic equivalence of microscopic and coarse-grained systems is ensured. The fluctuating dynamics of coarse-grained states generalize Macroscopic Fluctuation Theory (MFT) for non-reciprocal systems [88]. We obtain a non-quadratic dissipation function, which plays an important role in the exact quantification of the EPR, particularly for far-from-equilibrium systems [89, 90]. We decompose the EPR into its linearly independent contributions through an orthogonal decomposition, consisting of four components: relaxation of the reciprocal interaction energy functional, non-reciprocal interaction, coupling to the external reservoir, and external driving work. We show that non-reciprocal phase transitions are equivalent to dynamical phase transitions and are characterized by different scaling regimes for the non-reciprocal EPR, which physically correspond to sustaining dissipative vorticity currents.
@@
Finally, we demonstrate non-reciprocal and non-equilibrium analogs of reciprocal systems. For instance, irreversible thermodynamic relations: Onsager’s non-reciprocal relations, fluctuation-response relations, higher-order non-linear responses, and stochastic thermodynamic relations: fluctuation relations, thermodynamic kinetic uncertainty relations, and information thermodynamics, which are generalized to macroscopic systems. In the deterministic macroscopic limit, we formulate irreversible thermodynamics, Arrhenius transition state theory (LDB), and reaction-diffusion systems (dynamics) for interacting non-reciprocal density fields [91, 64, 65]. We elaborate on the application of our framework using a few prototypical models. Overall, our framework extends the applicability of stochastic thermodynamics for non-reciprocal and externally driven systems across different observation scales.
@@ 2 Microscopic Description
We consider a multi-particle system that comprises several particle types and resides in a discretized space: a regular lattice with lattice constant $l$, continuous time, and periodic boundary conditions. Physically, the lattice constant corresponds to using the microscopic diffusive length scale of the particles as the unit of spatial measurement. The continuum limit $l\to 0$ leads to an oversimplified coarse-grained description that does not account for the system’s inherent diffusive lengthscale. Therefore, we set $l=1$ throughout this paper unless we explicitly specify the continuum space limit $l\to 0$. The number of particles of type $i$ at lattice site index #\# is denoted by $N_{i}^{\#}$. The configuration of the lattice is denoted by $\{N\}$, whose dimension is determined by the lattice size and the number of particle types. The volume of the $d$-dimensional lattice space is denoted by $\mathcal{V}$. The faster degrees of freedom form the environment, which provides the thermodynamic cost for all microstate changes of the system, consisting of the slower degrees of freedom.
@@ 2.1 Microscopic Boltzmann weight
A Theoretical formulation
The microscopic interaction energy experienced by the type $i$ particle due to the type $j$ particle is denoted by $v_{ij}$. $v_{ij}<0$ implies that the type $i$ particle is attracted towards the type $j$ particle, while $v_{ij}>0$ indicates a repulsive interaction. Analogously, $v_{ij}$ represents the thermodynamic cost, supported by the environment, to place a type $i$ particle in the presence of a type $j$ particle, see fig. 2(a). The violation of the actio=reactio symmetry implies $v_{ij}\neq v_{ji}$ microscopically. $v_{ij}=v_{ji}$ indicates that the particles are reciprocal and respect the actio=reactio symmetry. To separate and quantify the actio=reactio symmetry preserving and breaking components, we define the reciprocal $v_{ij}^{r}=(v_{ij}+v_{ji})/2$ and non-reciprocal $v_{ij}^{nr}=(v_{ij}-v_{ji})/2$ interaction energies. $v_{ij}^{r}$ and $v_{ij}^{nr}$ correspond to the symmetry-preserving and symmetry-breaking parts, respectively, which fixes ‘the orthogonal reference gauge’ such that $v_{ij}^{r}=v_{ji}^{r}$ and $v_{ij}^{nr}=-v_{ji}^{nr}$ by construction.
The reciprocal and non-reciprocal components of the microscopic Boltzmann weight $\epsilon_{i}^{\#}$ of the type $i$ particle at #\# are denoted as $\epsilon_{i}^{r}$ and $f_{i}^{nr}$, respectively  [^2] Despite the equivalence of $\epsilon_{i}^{\#}$ to a chemical potential, we avoid calling so. Rather, it appears in eq. 3 as the argument in the exponential defining the transition rates for a single particle. If all particles are reciprocal, the usual physical meaning of the chemical potential is recovered. However, in principle, what we refer to as the ‘Boltzmann weight’ throughout this work is mathematically and physically equivalent to the ‘chemical potential’, but defined for a non-reciprocal and far-from-equilibrium system. This remark clarifies the nomenclature. , such that $\epsilon_{i}^{\#}=\epsilon_{i}^{r}+f_{i}^{nr}$,